OpenStudy (anonymous):

A hotel lobby can seat 14 people at chairs and loveseats. The lobby has a total of 11 pieces of furniture for seating. Assuming a chair seats 1 person and a loveseat seats 2, how many loveseats are in the lobby?

OpenStudy (anonymous):

14 = 1(l) + 1(11 - l) 14 = 2l + 12 - 1l 14 - 11 = 2l - 1l 3 = 1l l = 3 The Answer is 3.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

thats not right